OpenStudy (anonymous):

Find the zeros of the polynomial function. f(x) = x2 + 12x + 20 A. 2, 10 B. –4, –5 C. –2, –10 D. 4, 5

OpenStudy (***[isuru]***):

hey, this one is same as before... the question is asking u to solve it for "x" when 0 = x2 + 12x + 20 so... factorize it.. x^2 + 12x +20 =0 x^2 + 10x +2x +20 = 0 x(x + 10) +2(x +10) =0 (x + 2)(x + 10) = 0 which means.. x = -2 or x = -10 did ya get it ?

OpenStudy (anonymous):

I kinda get it but polynomial function is very hard for so.....but thanks for helping :D